OpenStudy (deadshot):

If f(x) = x plus 2, all over 2, then what is f-1(x) ?

OpenStudy (anonymous):

Switch the variables. Call f(x) y. So y = (x+2)/2 becomes x = (y+2)/2. Solve for y. Here, you get y = 2x - 2.

pooja195 (pooja195):

then the inverse is x = (f(x)+2)/2 solve for f(x) 2x = f(x) + 2 f(x) = 2x - 2 so u get 2x – 2

OpenStudy (deadshot):

okay, thanks!